Can you please write it down 350 words . I appreciate your help. Type it please I can’t read most of the handwriting.

Discussion Time Greg wants to help provide for the human rights and dignity of those left homeless due to Katrina. He has gathered a group of volunteers to go to meet the stream of humanity that is coming out of the worst hit area in search of food and shelter. However, how can he offer so many people help when shelter, food, and water are in short supply? How can you help Greg with this humanitarian effort? What would you do? all ows each person to start one discussion t scussion topic Forum Gradebook Started by Unread Francis Sanders
Answer #1

Solution:-

In terms of crisis or any emergency situations like in Katrina, all the volunteers should provide the necessary help to the people suffering in various forms. They ought to give them sustenance, protect, restorative offices and so on upto the degree conceivable. The sustenance or safe house might be hard to come by however circumstance requests sharing of everything that can be accessible with the goal that each individual enduring can get by in this circumstance.

We as a volunteer would help Greg with this compassionate exertion in each conceivable way. we should meet the general population who had endured because of this and give them nourishment, protect and different offices as indicated by their necessities. We ought to likewise guarantees and attempt to influence them to overlook that terrible a great time and endeavor to expedite grin their countenances.

Recall the dice game played between Alice and Bob: Alice rolls two dice, and tells Bob the sum; then Bob guesses a number from 1 to 6. Bob wins the game if his guess appears on either of the two dice. Now suppose Alice and Bob play for money. On the off chance that Bob wins, he wins a dollar, and on the off chance that he loses, Alice wins a dollar. We can demonstrate this utilizing an arbitrary variable W speaking to Bob's rewards. We saw that utilizing an ideal methodology, Bob wins with likelihood 5/9. This implies E[W] = (1)(5/9) + (−1)(4/9) = 1/9. To make the amusement all the more intriguing, Alice enables Bob to "twofold down": this implies after Alice discloses to Bob the entirety, Bob is permitted to twofold his wagered from one dollar to two dollars, in the event that he so picks. Give an ideal technique for Bob (for the two his figure and the twofold down choice) and register E[W] for this methodology.

if Z is a random variable representing the sum of the two dice, compute E[W] using the law of total expectation: E[W] = X 12 `=2 E[W | Z = `] Pr[Z = `].

Please do liked me..............